OpenStudy (anonymous):

Anyone good at Adv. Algebra w/ Financial Applications??? :) Ken and Kim have obtained a 30-year, fixed rate mortgage for $625,250 with a 7.05% interest rate. They purchased 3 points and their rate is now 6.775%. Factoring in the cost of points, when is the break-even point on their mortgage? I know how to find the break-even point. I just need to know if I find the monthly payment using the first given interest rate or find the difference between that monthly payment and the other monthly payment using the second given interest rate. :/

OpenStudy (dumbcow):

For loans involving monthly payments, the structure is a geometric series present value equation. You could also use a financial calculator. Since this is a math class i assume you have to find monthly payment by hand? Here is formula: \[PV = M (\frac{1-v^n}{1-v})\] \[v = \frac{1}{1+\frac{i}{12}}\]

OpenStudy (anonymous):

I have to use the formula M= i(1+i)^nt/(1+i)^nt-1 to calculate the monthly payment. I did this already. I just don't know whether I need to calculate the monthly payment using just the first given interest rate or calculate the monthly payment using both given interest rates.

OpenStudy (dumbcow):

calculate using both then look at difference in payment....the break even point is the month where the extra money paid (the difference) equals the extra initial payment in points (3% of loan)

OpenStudy (dumbcow):

I get a break even point of 163 months Dont buy the points unless you plan on staying in house for more than 13 years

OpenStudy (anonymous):

I did, too! I converted it to years and months. I got approximately 13 years and 7 months. Thank you for helping me. :)

OpenStudy (dumbcow):

yw